1. 9(x-5)-x(x-5)=
2. ax+ay+2bx+2by=
3. ma+mb+na+nb=
Can you solve and tell me how to do these three problems on factoring?

Found 2 solutions by Alex, gsmani_iyer:
Answer by Alex(3)

1. Take out the common factor (in this case, x-5)
(9-x)(x-5)
2. Factorise the first two terms and the second two terms:
a(x+y)+2b(x+y)
Then take out the common factor (as in question 1)
(a+2b)(x+y)
3. Factorise the first two terms and the second two terms:
m(a+b)+n(a+b)
Then take out the common factor (as in questions 1 & 2)
(m+n)(a+b)

Answer by gsmani_iyer(201)

1. 9(x-5)-x(x-5)
= (x-5)(9-x) Answer.
2 ax+ay+2bx+2by
= a(x+y)+2b(x+y)
= (x+y)(a-2b) Ans.
3. ma+mb+na+nb
= m(a+b)+n(a+b)
= (a+b)(m+n) Ans.
Try to take out the common term first. In the first problem, (x-5) is the common term. Take the common term as one factor and take the coefficients as the other factor. Here, in this case, 9 and -x are the coefficients. Take it as it is in order which comes to (9-x). That's all.
In the 2nd & 3rd problems, split out the expression into 2 terms each in such a way that in each two terms you have a common variable and factorise them.
Then again, take out the common term as one factor and the coefficients as another.
I hope you understood the basic principle of solving such problems. O.K